Grade 12th passGeneral Physics

if the velocity of a particle is given by 10sin(100πt+π/4) ,calculate its acceleration at t=0s

Marib
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear student
v = 10 sin(100 \pit + \pi/4)
On differentiation
a = dv/dt = 10 cos(100\pit + \pi/4) * 100\pi
Now at t = 0
a = 1000\pi * cos(\pi/4)
Regards
Arun (askIITians forum expert)